I need to solve this by substitution or addition, whichever one is easier.
2y-x=3
x=3y-5

Answer by KnightOwlTutor(293)

2y-x=3
x=3y-5
Replace the x in the first equation with 3y-5
2y-3y+5=3
Simplify the y term
-y+5=3
Subtract 5 from both sides
-y=-2
Divide both sides by -1
y=2
x=6-5
x=1

2y-x=3
x=3y-5
Let's plug in the calculated x and y values to confirm
4-1=3
1=6-5